This happened to us once before. the trapped hovered close to zero instead of our usual ~0.5 and box was close to 1.5. signal intensity was normal but the delta values were very noisy and probably the beam was not stable. the solution was to open the source again, detach the filament and then reattach it.
Hagit P. Affek
Dept. of Geology & Geophysics
210 Whitney Ave.
PO Box 208109
New Haven, CT 06520-8109
On 12/19/14, 11:11 AM, Herbert Tobias wrote:
This week I cleaned the source of our MAT 253 and installed a new Thermo filament. Apart from resolving a bunch of annoying issues over the days, and re-cleaning yesterday, the Trap current on the system is currently reading Zero (0), while the Box current reads the emission I set in the focus settings (i.e. 1.5 mA or 1.0 mA etc). All through the week and just this morning the Box and Trap current was reading approximately equal (i.e. 0.75 mA and 0.75 mA). Right now for the past couple hours it is lopsided as noted above (Box= 1.5 mA and Trap = 0 mA). But the instrument does provide strong signal for CO2 (m/z 44 = ~4500 mV at 1.5 bar on Ref 1 of a Conflo III). I measure no shorts to the Trap. I decided to start baking the system for awhile and see what happens. Any thoughts?
(Title of e-mail sounds like a psychological thriller, eh?)
thanks and regards
Dr. Herbert Tobias
Sr. Research Associate
Division of Nutritional Sciences
244 Garden Ave
Ithaca NY 14853